OpenStudy (anonymous):

DERIVATE PLEASE!! :D y = (cot ^2 x)/(1+x^2)

OpenStudy (solomonzelman):

ok, at first what is the derivative of cotx?

OpenStudy (solomonzelman):

(I am assuming that by "derivate" you mean differentiate.)

OpenStudy (solomonzelman):

So you need to find: \(\large\color{black}{ \displaystyle \frac{ d}{dx} \left[ \cot~x\right] =~\color{red}{\text{____}}}\) and then cot^2(x) goes by the chain rule, \(\large\color{black}{ \displaystyle \frac{ d}{dx} \left[ (\cot~x)^\color{blue}{2}\right] =\color{blue}{2}(\cot~x)^{\color{blue}{2}-1}~\times~\color{red}{\text{____}}}\)

OpenStudy (solomonzelman):

the red is the blank for the derivative of cot x.

OpenStudy (anonymous):

Is right?